OpenStudy (anonymous):

Simplify the expression, using scientific notation.(9 x 10^2)(2 x 10^10) 1.1x10^21 1.1x10^13 1.8x10^21 1.8x10^13

OpenStudy (cwrw238):

(9 x 10^2)(2 x 10^10) = 9*2 * (10^2 * 10^10) = 18 * 10^12 can u continue?

OpenStudy (anonymous):

no idea lost..

OpenStudy (cwrw238):

9 * 2 = 18 10^2 * 10^10 = 10^(10+2) = 10^12 you add the indices

OpenStudy (anonymous):

yeah i understand that but 1.1x10^21, 1.1x10^13, 1.8x10^21, 1.8x10^13, ar ethe only answers

OpenStudy (cwrw238):

ok scientific notation is defined as N.NN.. * 10^n

OpenStudy (cwrw238):

the first number must be between 1 and 10:

OpenStudy (cwrw238):